# Advection selects pattern in multi-stable emulsions of active droplets

### v1

@@ Abstract
Controlling the size of droplets, for example in biological cells, is challenging because large droplets typically outcompete smaller droplets due to surface tension. This coarsening is generally accelerated by hydrodynamic effects, but active chemical reactions can suppress it. We show that the interplay of these processes leads to three different dynamical regimes: (i) Advection dominates the coalescence of small droplets, (ii) diffusion leads to Ostwald ripening for intermediate sizes, and (iii) reactions finally suppress coarsening. Interestingly, a range of final droplet sizes is stable, of which one is selected depending on initial conditions. Our analysis demonstrates that hydrodynamic effects control initial droplet sizes, but they do not affect the later dynamics, in contrast to passive emulsions.
Emulsions of multiple droplets provide spatial structures in biological cells Brangwynne et al. (2009); Hyman et al. (2014); Banani et al. (2017); Zwicker et al. (2025) and synthetic applications Jambon-Puillet et al. (2024); Song et al. (2025). In these examples, effective patterning requires controlled droplet sizes and positions. These aspects are affected by physical processes such as diffusion, advection, and chemical reactions Zwicker et al. (2025). In particular, diffusion generally causes droplet coarsening, either because molecules diffuse from smaller to larger droplets (Ostwald ripening Ostwald (1897); Voorhees (1985)) or because droplet diffusion leads to coalescence Tanaka (1997). Hydrodynamic advection generally accelerates coarsening Nikolayev et al. (1996); Tanaka (1997); Huo et al. (2003), whereas driven chemical reactions can suppress coarsening Wurtz and Lee (2018); Bressloff (2020); Sastre et al. (2024); Bauermann et al. (2024, 2025). However, how the interplay of these processes organizes emulsions, and in particular selects length scales, is poorly understood.
@@
where the left hand side describes the material derivative with velocity field $\bm{v}$, whereas the right hand side accounts for diffusion with constant mobility $\Lambda$ and a linear reaction with rate $k$ describing the conversion between A and B; the net conversion vanishes at $c=c_{0}$. The diffusive flux is driven by gradients in the exchange chemical potential $\mu=\delta F/\delta c$, which is derived from a free energy $F$. As a model for phase separation, we choose the Ginzburg–Landau form Weber et al. (2019)
@@
where $a$ sets the energy scale, $c_{\mathrm{in}}$ denotes the concentration difference between coexisting phases, and $\kappa$ penalizes gradients, implying a surface tension of $\gamma=\frac{1}{6}c_{\mathrm{in}}^{2}\sqrt{\kappa\,a}$ and an interfacial width of $w=\sqrt{\kappa/a}$ Weber et al. (2019).
We consider small systems with low velocities, implying that inertial effects are negligible. Consequently, the velocity field $\bm{v}$ is governed by the Stokes equations,
@@
Figure 1: Turnover causes hexagonal droplet pattern. (A) Schematic of our model involving two species that inter-convert and phase separate from each. (B) Simulation snapshot showing $c(\bm{r})$ at stationarity for $\mathrm{P\hskip-1.0pte}=0$, $c_{0}=0.35\,c_{\mathrm{in}}$, $k=0.02/\tau$, $\tau=w^{2}/D$, and $w=\sqrt{\kappa/a}$.
@@
Figure 2: Reaction rate controls stationary pattern size. (A) Pattern size $L_{\mathrm{eq}}$, determined numerically (symbols) and predicted by Eq. (6) (lines), as a function of reaction rate $k$ for various average compositions $c_{0}$. Inset shows corresponding radii. (B) Concentration profile $c(\bm{r})$ for $k=10^{-3}/\tau$ in a unit cell with adjustable length $L$. (C) Average pattern size $\bar{L}$ as a function of time for various $k$, where $\bar{L}=2\pi^{-1/2}(V_{\mathrm{sys}}/N)^{1/2}$ from droplet count NN and system size $V_{\mathrm{sys}}$. Numerical data is compared to $L_{\mathrm{eq}}$ from Eq. (6) (gray dotted lines) and a $t^{1/3}$ power law (black dashed line). (D) Distribution of droplet radii $R$ (left) and the average radius $\bar{R}$ (right) for various time points for $k=10^{-5}/\tau$. Numerical data is compared to the prediction from spinodal decomposition (gray dashed line) and $R_{\mathrm{eq}}=\sqrt{c_{0}/4c_{\mathrm{in}}}\,L_{\mathrm{eq}}$ (black dashed line). (A–D) Model parameters are $c_{0}=0.25\,c_{\mathrm{in}}$, $\tau=w^{2}/D$, and $w=\sqrt{\kappa/a}$.
We start by investigating the interplay between diffusion and reactions, neglecting hydrodynamics for now ($\mathrm{P\hskip-1.0pte}=0$, implying $\bm{v}=\bm{0}$). Numerical simulations suggest that reactions suppress coarsening, so that droplets form a hexagonal pattern (Fig. 1B). To understand this, we map the active system described by Eq. (1) to an equivalent passive system using a surrogate free energy Liu and Goldenfeld (1989); Muratov (2002)
@@
This potential captures the effect of reactions as non-local interactions, so that the dynamics $\partial_{t}c=\Lambda\nabla^{2}\delta\tilde{F}/\delta c$ are identical to Eq. (1) when $\bm{v}=\bm{0}$. In particular, we minimize $\tilde{F}$ to identify stationary states (Appendix B.2). The pattern size, quantified by the center-to-center distance $L$, decreases with higher reaction rates $k$ (Fig. 2A), consistent with previous work Glotzer et al. (1995); Christensen et al. (1996); Muratov (2002); Zwicker et al. (2015).
We next derive an analytical estimate of the pattern size $L$ by focusing on a single droplet of radius $R$ and approximate its environment as a spherically symmetric shell with radius $\frac{1}{2}L$ (Appendix B.3). To determine the relation between $R$ and $L$, we focus on two-dimensional systems and approximate the concentration profile as $c(r)=c_{\mathrm{in}}\Theta(R-r)$, which describes droplets without reactions and surface tension effects. Since the average concentration is $c_{0}$ in steady state, we conclude that $R/L=\sqrt{c_{0}/4c_{\mathrm{in}}}$, based on a spherically symmetric system. Inserting this together with the solution of Eq. (5) into Eq. (4) leads to an expression for $\tilde{F}$ that only depends on $L$. Minimizing $\tilde{F}$ with respect to $L$ yields (Appendix B.3)
@@
consistent with our data (Fig. 2A). This expression is proportional to Eq. (58) in ref. Muratov (2002), but our pre-factor approximates the data better. In any case, material turnover leads to stationary states with controlled droplet sizes with $L_{\mathrm{eq}},R_{\mathrm{eq}}\propto k^{-1/3}$.
We next analyze the approach toward stationary state. To analyze generic situations, we start with a slightly perturbed homogeneous state and simulate spinodal decomposition (Appendix A.1). Once proper droplets with a narrow size distribution form, we switch to an effective simulation method to cover long time scales Kulkarni et al. (2023). The average pattern size $\bar{L}$ generally increases over time and saturates at a level that depends on $k$ (Fig. 2C). During the coarsening phase, we find $\bar{L}\propto t^{1/3}$ (Fig. 2C), consistent with Lifshitz–Slyzov–Wagner theory Lifshitz and Slyozov (1961); Wagner (1961). Moreover, our data suggests that the droplet size distribution is similar to the expected universal one during coarsening Lifshitz and Slyozov (1961); Wagner (1961), but it becomes much narrower close to the stationary state (Fig. 2D) Bauermann et al. (2025). Interestingly, the pattern size saturates below the length scale predicted by Eq. (6) (Fig. 2C).
@@
To understand the minimal size of patterns, we next analyze their stability. Our simulations indicate that patterns coarsen by dissolving droplets, whereas coalescence is negligible in the late stage of coarsening. To determine the smallest droplet that can be stable, we consider a system comprising a droplet of radius $R_{1}$, surrounded by droplets of radii $R_{2}$ at distance $L$, assuming spherical symmetry (Fig. S3). The dynamics of $R_{1}$ are governed by $\partial_{t}R_{1}=(j_{\mathrm{in}}-j_{\mathrm{out}})/c_{\mathrm{in}}$, where $j_{\mathrm{in}}$ and $j_{\mathrm{out}}$ are the diffusive fluxes inside and outside of the interface Weber et al. (2019). These fluxes can be determined analytically assuming quasi-stationarity, implying that $\partial_{t}R_{1}$ can be expressed as a function of $R_{1}$, $R_{2}$, and $L$. A stability analysis for small perturbations around $R_{1}=R_{2}=R_{*}$ allows us to determine the minimal stable radius, $R_{\mathrm{min}}$ (Appendix C). The corresponding predicted minimal pattern size $L_{\mathrm{min}}$ underestimates the smallest observed patterns (Fig. 3B), but it reveals the correct trends: Droplets can be smaller if there is more material (larger $c_{0}$). The dependence of $L_{\mathrm{min}}$ on $c_{0}$ is stronger than for $L_{\mathrm{eq}}$, but both quantities scale as $k^{-1/3}$ (Fig. S4B and Zwicker et al. (2015)). The stability analysis predicts that patterns with droplets above the minimal size $R_{\mathrm{min}}$ are stable.
@@
Figure 4: Advection accelerates coarsening. Pattern size $\bar{L}$ (shaded area indicates SD, $n=20$) as a function of time $t$ for various Péclet numbers $\mathrm{P\hskip-1.0pte}=w^{2}\,a\,c_{\mathrm{in}}^{2}/(D\eta)$ and reaction rates $k$ (across panels). Black dashed lines show power laws $\bar{L}\propto t$ (left) and $\bar{L}\propto t^{1/3}$ (right). Gray dotted lines show $L_{\mathrm{eq}}$ calculated from Eq. (6). Inset in (B) shows the velocity field generated by two colliding droplets (gray arrows) for $k=10^{-4}/\tau$ and $\mathrm{P\hskip-1.0pte}=100$. (D–F) Volume averaged viscous (blue) and diffusive (orange) dissipation corresponding to (A–C) for $\mathrm{P\hskip-1.0pte}=100$ (Appendix E). The curves were smoothed by convolution with a Gaussian kernel with a standard deviation of 2 data-points to make the plot more readable. Black dashed line in (D) shows a power law $\propto t^{-2/3}$. (A–F) Model parameters are $c_{0}=0.4\,c_{\mathrm{in}}$, $\tau=w^{2}/D$, and $w=\sqrt{\kappa/a}$.
Finally, we ask how hydrodynamics affects the coarsening of chemically active droplets by considering $\mathrm{P\hskip-1.0pte}>0$. We perform numerical simulations of Eqs. (1)–(3), where we solve the Stokes equation using a pseudo-spectral method (Appendix A). To form droplets effectively, we consider the spinodal decomposition regime. We find that hydrodynamics generally accelerates coarsening for all reaction rates (Fig. 4), consistent with literature Nikolayev et al. (1996); Tanaka (1997); Huo et al. (2003). This acceleration is less pronounced for a smaller average fraction $c_{0}$ (Fig. S1), indicating that it relies on a large droplet density.
For passive mixtures ($k=0$, Fig. 4A), our data at early times is consistent with the theoretically expected scaling of $\bar{L}\propto t$ resulting from “coalescence-induced coalescence” in the viscous regime Nikolayev et al. (1996); Tanaka (1997); Wagner and Cates (2001). In contrast, we observe a transition to Ostwald ripening at large length and time scales, where the mean pattern size obeys $\bar{L}\propto t^{1/3}$. Interestingly, we observe a plateau region between these two regimes where $\bar{L}$ hardly changes, suggesting that Ostwald ripening is delayed. Note that the transition between the two regimes is opposite to what has been described for bicontinuous structures in equivalent systems, where diffusion dominates at small scales (implying $\bar{L}\propto t^{1/3}$) and advection becomes important later (leading to $\bar{L}\propto t$Siggia (1979). This difference might be caused by the fact that advection is less important in our system as judged by the smaller viscous dissipation compared to diffusive dissipation (Fig. 4D). However, the identical scaling of the dissipations indicates that both advection and diffusion are relevant for the entire coarsening process of passive mixtures.

### v2

# Advection selects pattern in multistable emulsions of active droplets

@@ Abstract
Controlling the size of droplets, for example in biological cells, is challenging because large droplets typically outcompete smaller droplets due to surface tension. This coarsening is generally accelerated by hydrodynamic effects, but active chemical reactions can suppress it. We show that the interplay of these processes leads to three different dynamical regimes: (1) Advection dominates the coalescence of small droplets, (2) diffusion leads to Ostwald ripening for intermediate sizes, and (3) reactions finally suppress coarsening. Interestingly, a range of final droplet sizes is stable, of which one is selected depending on initial conditions. Our analysis demonstrates that hydrodynamic effects control initial droplet sizes, but they do not affect the later dynamics, in contrast to passive emulsions.
I Introduction
Emulsions of multiple droplets provide spatial structures in biological cells Brangwynne et al. (2009); Hyman et al. (2014); Banani et al. (2017); Zwicker et al. (2025) and synthetic applications Jambon-Puillet et al. (2024); Song et al. (2025). In these examples, effective patterning requires controlled droplet sizes and positions. These aspects are affected by physical processes such as diffusion, advection, and chemical reactions Zwicker et al. (2025). In particular, diffusion generally causes droplet coarsening, either because molecules diffuse from smaller to larger droplets (Ostwald ripening Ostwald (1897); Voorhees (1985)) or because droplet diffusion leads to coalescence Tanaka (1997). Hydrodynamic advection generally accelerates coarsening Nikolayev et al. (1996); Tanaka (1997); Huo et al. (2003), whereas driven chemical reactions can suppress coarsening Wurtz and Lee (2018); Bressloff (2020); Sastre et al. (2024); Bauermann et al. (2025a, b). However, how the interplay of these processes organizes emulsions, and in particular selects length scales, is poorly understood.
II Results
II.1 Chemically active droplets couple diffusion, advection, and reactions
@@
where the left-hand side describes the material derivative with velocity field $\bm{v}$, whereas the right-hand side accounts for diffusion with constant mobility $\Lambda$ and a linear reaction with rate $k$ describing the conversion between A and B. This conversion reaction describes a driven reaction since it is not derived based on thermodynamic principles Lefever et al. (1995); Kirschbaum and Zwicker (2021); Zwicker et al. (2025). In particular, the activity-dependent reaction equilibrium at $c=c_{0}$ is distinct from the thermodynamic equilibrium described by the free energy $F$ of the system. In contrast, the diffusive flux is driven by gradients in the exchange chemical potential $\mu=\delta F/\delta c$, implying that it biases the system toward the free energy minimum. To analyze the interplay of passive diffusion and active reactions, we consider a free energy of the Ginzburg–Landau form Weber et al. (2019)
@@
where $a$ sets the energy scale, $c_{\mathrm{in}}$ denotes the concentration difference between coexisting phases, and $\kappa$ penalizes gradients, implying a surface tension of $\gamma=\frac{1}{6}c_{\mathrm{in}}^{2}\sqrt{\kappa\,a}$ and an interfacial width of $w=\sqrt{\kappa/a}$ Weber et al. (2019). The concentrations of coexisting equilibrium phases are thus $c=0$ and $c=c_{\mathrm{in}}$, which are distinct from the reaction equilibrium at $c=c_{0}$ since the reactions are active. This imbalance drives continuous fluxes and flows in the system, which we analyze in small systems with low velocities, implying that inertial effects are negligible. Consequently, the velocity field $\bm{v}$ is governed by the Stokes equations,
@@
Figure 1: Turnover causes hexagonal droplet pattern. (A) Schematic of our model involving two species that inter-convert and phase separate from each other. (B) Simulation snapshot showing $c(\bm{r})$ at stationarity for $\mathrm{P\hskip-1.0pte}=0$, $c_{0}=0.35\,c_{\mathrm{in}}$, $k=0.02/\tau$, $\tau=w^{2}/D$, and $w=\sqrt{\kappa/a}$.
@@
II.2 Chemical reactions arrest droplet coarsening
Figure 2: Reaction rate controls stationary pattern size. (A) Pattern size $L_{\mathrm{eq}}$, determined numerically (symbols) and predicted by Eq. (6) (lines), as a function of reaction rate $k$ for various average compositions $c_{0}$. Inset shows corresponding radii. (B) Concentration profile $c(\bm{r})$ for $k=10^{-3}/\tau$ in a unit cell with adjustable length $L$. (C) Average pattern size $\bar{L}$ as a function of time for various $k$, where $\bar{L}=2\pi^{-1/2}(V_{\mathrm{sys}}/N)^{1/2}$ from droplet count NN and system size $V_{\mathrm{sys}}$. Numerical data are compared to $L_{\mathrm{eq}}$ from Eq. (6) (gray dotted lines) and a $t^{1/3}$ power law (black dashed line). (D) Distribution of droplet radii $R$ (left) and the average radius $\bar{R}$ (right) for various time points for $k=10^{-5}/\tau$. Numerical data are compared to the prediction from spinodal decomposition (gray dashed line) and $R_{\mathrm{eq}}=\sqrt{c_{0}/4c_{\mathrm{in}}}\,L_{\mathrm{eq}}$ (black dashed line). (A–D) Model parameters are $c_{0}=0.25\,c_{\mathrm{in}}$, $\tau=w^{2}/D$, and $w=\sqrt{\kappa/a}$.
II.2.1 Stationary droplets are larger for weaker reactions
We start by investigating the interplay between diffusion and reactions, neglecting hydrodynamics for now ($\mathrm{P\hskip-1.0pte}=0$, implying $\bm{v}=\bm{0}$). Numerical simulations in two dimensions suggest that reactions suppress coarsening, so that droplets form a hexagonal pattern (Fig. 1B). To understand how chemical reactions alter material fluxes to suppress coarsening, we map the active system described by Eq. (1) to an equivalent passive system with non-local interactions. In particular, we rewrite the dynamics using a surrogate free energy Liu and Goldenfeld (1989); Muratov (2002)
@@ II.2.1 Stationary droplets are larger for weaker reactions
This potential captures the effect of reactions as nonlocal interactions, so that the dynamics $\partial_{t}c=\Lambda\nabla^{2}\delta\tilde{F}/\delta c$ are identical to Eq. (1) when $\bm{v}=\bm{0}$. Eqs. (4)–(5) reveal that the reactions are analogous to an electrostatics problem, where $(c-c_{0})$ acts as a charge density creating an electrostatic potential $\psi$. Droplets can thus be interpreted as positively charged disks, surrounded by a negative charge cloud extending over the reaction-diffusion length $\xi=\sqrt{D/k}$ Zwicker et al. (2025); Ziethen and Zwicker (2024). If the average concentration of the system reached its stationary state value $c_{0}$, the mathematical mapping to the equilibrium system exhibits the same dynamics as the original system given by Eq. (1). Consequently, minimizing $\tilde{F}$ allows us to identify stationary states of the original system. Our numerical minimization described in Appendix B.2, reveals that the pattern size, quantified by the center-to-center distance $L$, decreases with higher reaction rates $k$ (Fig. 2A), consistent with previous work Glotzer et al. (1995); Christensen et al. (1996); Muratov (2002); Zwicker et al. (2015).
We next derive an analytical estimate of the pattern size $L$ by focusing on a single droplet of radius $R$ and approximating its environment as a spherically symmetric shell with radius $\frac{1}{2}L$ (Appendix B.3). To determine the relation between $R$ and $L$, we focus on two-dimensional systems and approximate the concentration profile as $c(r)=c_{\mathrm{in}}\Theta(R-r)$, which describes droplets without reactions and surface tension effects. Since the average concentration is $c_{0}$ in steady state, we conclude that $R/L=\sqrt{c_{0}/4c_{\mathrm{in}}}$, based on a spherically symmetric system. Inserting this together with the solution of Eq. (5) into Eq. (4) leads to an expression for $\tilde{F}$ that only depends on $L$. Minimizing $\tilde{F}$ with respect to $L$ yields (Appendix B.3)
@@
consistent with our data (Fig. 2A). This expression is proportional to Eq. (58) in ref. Muratov (2002), but our prefactor approximates the data better. In any case, material turnover leads to stationary states with controlled droplet sizes with $L_{\mathrm{eq}},R_{\mathrm{eq}}\propto k^{-1/3}$.
II.2.2 Chemically active droplets coarsen until their stationary size
We next analyze the approach toward stationary state. To analyze generic situations, we start with a slightly perturbed homogeneous state and simulate spinodal decomposition (Appendix A.1). Once proper droplets with a narrow size distribution form, we switch to an effective simulation method to cover long timescales Kulkarni et al. (2023). The average pattern size $\bar{L}$ generally increases over time and saturates at a level that depends on $k$ (Fig. 2C). During the coarsening phase, we find $\bar{L}\propto t^{1/3}$ (Fig. 2C), consistent with Lifshitz–Slyzov–Wagner theory Lifshitz and Slyozov (1961); Wagner (1961). Moreover, our data suggest that the droplet size distribution is similar to the expected universal one during coarsening Lifshitz and Slyozov (1961); Wagner (1961), but it becomes much narrower close to the stationary state (Fig. 2D) Bauermann et al. (2025b). Interestingly, the pattern size saturates below the length scale predicted by Eq. (6) (Fig. 2C).
II.2.3 Chemical reactions introduce multistability
@@
To understand the minimal size of patterns, we next analyze their stability. Our simulations indicate that patterns coarsen by dissolving droplets, whereas coalescence is negligible in the late stage of coarsening. To determine the smallest droplet that can be stable, we consider a system comprising a droplet of radius $R_{1}$, surrounded by droplets of radii $R_{2}$ at distance $L$, assuming spherical symmetry (Fig. S3). The dynamics of $R_{1}$ are governed by $\partial_{t}R_{1}=(j_{\mathrm{in}}-j_{\mathrm{out}})/c_{\mathrm{in}}$, where $j_{\mathrm{in}}$ and $j_{\mathrm{out}}$ are the diffusive fluxes inside and outside of the interface Weber et al. (2019). These fluxes can be determined analytically assuming quasistationarity, implying that $\partial_{t}R_{1}$ can be expressed as a function of $R_{1}$, $R_{2}$, and $L$. A stability analysis for small perturbations around $R_{1}=R_{2}=R_{*}$ allows us to determine the minimal stable radius, $R_{\mathrm{min}}$ (Appendix C). The corresponding predicted minimal pattern size $L_{\mathrm{min}}$ underestimates the smallest observed patterns (Fig. 3B), but it reveals the correct trends: Droplets can be smaller if there is more material (larger $c_{0}$). The dependence of $L_{\mathrm{min}}$ on $c_{0}$ is stronger than for $L_{\mathrm{eq}}$, but both quantities scale as $k^{-1/3}$ (Fig. S4 and Zwicker et al. (2015)). The stability analysis predicts that patterns with droplets above the minimal size $R_{\mathrm{min}}$ are stable.
@@
II.3 Hydrodynamic advection accelerates coarsening by inducing coalescence
Figure 4: Advection accelerates coarsening. Pattern size $\bar{L}$ (shaded area indicates standard deviation, $n=20$) as a function of time $t$ for various Péclet numbers $\mathrm{P\hskip-1.0pte}=w^{2}\,a\,c_{\mathrm{in}}^{2}/(D\eta)$ and reaction rates $k$ (across panels). Black dashed lines show power laws $\bar{L}\propto t$ (left) and $\bar{L}\propto t^{1/3}$ (right). Gray dotted lines show $L_{\mathrm{eq}}$ calculated from Eq. (6). Inset in (B) shows the velocity field generated by two colliding droplets (gray arrows) for $k=10^{-4}/\tau$ and $\mathrm{P\hskip-1.0pte}=100$. (D–F) Volume averaged viscous (blue) and diffusive (orange) dissipation corresponding to (A–C) for $\mathrm{P\hskip-1.0pte}=100$ (Appendix E). The curves were smoothed by convolution with a Gaussian kernel with a standard deviation of 2 data-points to make the plot more readable. Black dashed line in (D) shows a power law $\propto t^{-2/3}$. (A–F) Model parameters are $c_{0}=0.4\,c_{\mathrm{in}}$, $\tau=w^{2}/D$, and $w=\sqrt{\kappa/a}$.
Finally, we ask how hydrodynamics affects the coarsening of chemically active droplets by considering $\mathrm{P\hskip-1.0pte}>0$. We perform numerical simulations of Eqs. (1)–(3), where we solve the Stokes equation using a pseudospectral method (Appendix A). To form droplets effectively, we consider the spinodal decomposition regime. We find that hydrodynamics generally accelerates coarsening for all reaction rates (Fig. 4), consistent with literature Nikolayev et al. (1996); Tanaka (1997); Huo et al. (2003). This acceleration is less pronounced for a smaller average fraction $c_{0}$ (Fig. S1), indicating that it relies on a large droplet density.
For passive mixtures ($k=0$, Fig. 4A), our data at early times is consistent with the theoretically expected scaling of $\bar{L}\propto t$ resulting from “coalescence-induced coalescence” in the viscous regime Nikolayev et al. (1996); Tanaka (1997); Wagner and Cates (2001); Gsell and Merkel (2022). In contrast, we observe a transition to Ostwald ripening at large length scales and timescales, where the mean pattern size obeys $\bar{L}\propto t^{1/3}$. Interestingly, we observe a plateau region between these two regimes where $\bar{L}$ hardly changes, suggesting that Ostwald ripening is delayed. Note that the transition between the two regimes is opposite to what has been described for bicontinuous structures in equivalent systems, where diffusion dominates at small scales (implying $\bar{L}\propto t^{1/3}$) and advection becomes important later (leading to $\bar{L}\propto t$Siggia (1979). This difference might be caused by the fact that advection is less important in our system as judged by the smaller viscous dissipation compared to diffusive dissipation (Fig. 4D). However, the identical scaling of the dissipations indicates that both advection and diffusion are relevant for the entire coarsening process of passive mixtures.